A local TV repairs shop uses 36,000 units of a part each year (A maximum consumption of 100 units per working day). It costs Rs. 20 to place and receive an order. The shop orders in lots of 400 units. It cost Rs. 4 to carry one unit per year of inventory.

 

Requirements:

(1) Calculate total annual ordering cost

(2) Calculate total annual carrying cost

(3) Calculate total annual inventory cost

(4) Calculate the Economic Order Quantity

(5) Calculate the total annual cost inventory cost using EOQ inventory Policy

(6) How much save using EOQ

(7) Compute ordering point assuming the lead time is 3 days




1
Expert's answer
2022-02-09T10:52:56-0500

1. Calculate total annual ordering cost

Ordering cost = fixed ordering cost (F) × Number of orders per year N

Ordering cost = 20 × "\dfrac{36000}{400}"

Ordering cost = Rs. 1800

2. Calculate total annual carrying cost

               Carrying cost = carrying cost (C) ×"\dfrac{EOQ}{2}"

               Carrying cost = 4×"\dfrac{400}{2}"

               Carrying Cost = Rs. 800

3. Calculate total annual inventory cost

               Total inventory Cost = Ordering cost + Carrying cost

               Total inventory cost = 1800 + 800

               Total inventory coost = Rs. 2600

4. Calculate the economic order quantity

               EOQ = "\sqrt{\smash[b]{\frac{2×RU×OC}{UC×CC\%}}}"

= "\sqrt{\smash[b]{\frac{2×36000×20}{4}}}"

EOQ = 600 units

5. Calculate the total annual cost inventory cost using EOQ inventory Policy

Total Inventory Cost = [fixed ordering cost (F) × Number of order per year N] + Carrying cost (C) × "\dfrac{EOQ}{2}"

                                               = [20 ×"\dfrac{36000}{600}]+[4 × \dfrac{600}{2}]"

                                               = 1200 + 1200

Total Inventory Cost = Rs. 2400

6. How much save using EOQ:

               Savings = 2600 – 2400

               Savings = Rs. 200

7. Compute ordering point assuming the lead time is 3 days.

Re-order level = maximum consumption × lead time[maximum]

Re-order level = 100 × 3

Re-order level = 300 per days